# Evaluating the Instruction-following Abilities of Language Models using Knowledge Tasks

### v1

@@ 4.3.1 Printing the correct answer
We begin our experiments with the simplest task – given a multiple-choice question with option labels and their texts, we instruct the models to print the text associated with the correct answer instead of the answer label. From a knowledge perspective, this task is no harder than selecting the answer label. However, as shown on the left in Figure 1, we observe a significant drop ($\sim 20\%$ on average) in knowledge-task performance when instructing the model to respond with the text associated with the answer instead of its label. The pattern is consistent for frontier models like GPT-4o on the Lite Benchmark (Figure 6).
(a) Small Scale Models: Full Benchmark
(b) Medium Scale Models: Full Benchmark
Figure 2: Performance variation (strict and loose) of exact match scores across the different answer-conditioned instruction categories (All) from Table 2 compared to corresponding performance on $print\_correct\_answer$ (PCA).
(a) Large Scale Models: Full Benchmark
(b) Frontier Models: Lite Benchmark
Figure 3: Performance variation (strict and loose) of exact match scores across the different answer-conditioned instruction categories (All) from Table 2 compared to corresponding performance on $print\_correct\_answer$ (PCA).
We hypothesize that this drop in performance could, in part, be due to the training process resulting in models being over-fit to certain input/output task formats, resulting in worse instruction following for other formats. Some common issues we observed include models outright ignoring the instruction and continuing to generate labels, or generating only Chain-of-Thought reasoning without a final answer, missing the output keyword specified in the prompt, etc, reflected by the knowledge and instruction following errors in Figure 1 on the right.
Inspired by this, and to further illustrate the challenges that LLMs face on simple instruction-following, we study their performance when the final output requires first inferring the correct answer, and then applying operations specified by the instructions on the correct answer.
@@ 4.3.2 Answer-conditioned instruction performance
We present results from different model scales across our five instruction categories in Figures 2(a),2(b),3(a),3(b). We compare this to their corresponding performance on the baseline task of $print\_correct\_answer$ (PCA).
Small-Scale and Medium-Scale models: We observe a $10\%-40\%$ drop in performance compared to the baseline across all instruction categories (Figures . 2(a) and 2(b)). In particular, we notice that all models struggle on our set of string and numeric manipulation instructions, suggesting a bias towards certain input/output format instructions (see Figure 25(a) and 25(b) for instruction-specific results for each model).
We further notice that models such as Qwen-2.5-14B and Llama-3.1-8B exhibit good loose evaluation scores for numeric manipulations and formatting the correct answer, but suffer a large drop in the corresponding strict evaluation. This difference suggests that these models are able to grasp the expectations from the instructions, but fail to follow them precisely. Examples of this include incrementing an answer by $0.1$ when asked to increment by one, or only returning the special string when instructed to append it to the correct answer text, or even adding/missing characters from the provided options when instructed to return them as is. Finally, all models also find the operations on list categories to be challenging – where interestingly the performance of models across both sets - conditional on correct answers vs. not, is similar.
Large-scale and Frontier models: The improved capabilities of these larger models are evident from the absolute improvement in performance as shown in Figures 3(a) and 3(b). We also observe a smaller drop in performance between their loose evaluation and strict evaluation scores, reflecting more precise instruction following. However, the trend of performance deterioration ($5\%-40\%$) across instructions, compared to their respective baseline knowledge-answering tasks, still persists in these models, demonstrating opportunities to improve their instruction following.
(a) Full Benchmark
(b) Lite Benchmark
Figure 4: Impact of distractor instructions on exact match performance across tasks and instructions, compared to its corresponding $print\_correct\_answer$ performance. A drop indicates the model getting distracted by an inapplicable instruction.
4.3.3 Instructions as distractors
@@
Our dataset also includes instructions that apply only when certain properties of a knowledge-task answer are fulfilled. For instance, instructions for incrementing the correct answer by one if numeric, formatting numeric values, and printing any numeric answers in words, do not apply on tasks with textual answers. They serve as distractors, and we expect model performance to be unaffected since these instructions are not applicable and do not alter the original knowledge-task answer.
However, from Figure 4, we observe that there is a 5-20% drop in small, medium, large, and frontier scale models. While model failures in the presence of distractors have been studied before (Shi et al., 2023; Feng et al., 2024), to the best of our knowledge this is the first work to study them in an instruction-following setting.
4.3.4 Knowledge-task characteristics and Instruction-following
(a) Medium Scale Models: Numeric Manipulation
(b) Frontier Models: Numeric Manipulation
Figure 5: Performance variation (strict and loose) of exact match scores for the Numeric Manipulation instruction category compared to its corresponding performance on $print\_correct\_answer$ (PCA).
As seen in Figure 5, the performance drop for models for an instruction category can also be dependent on the nature of the knowledge-task. For instance, models appear to have a larger relative drop on MathQA as compared to MMLUPro for the numeric manipulation instruction category. Models also struggle more on string manipulation operations on PIQA - probably because of the long sentences that are part of answer candidates. Other knowledge-task and model scales have been presented in Appendix Section A.5.
@@ 4.4 Benchmark
Table 4: Performance of the Medium, Large and Frontier Models on our Lite Benchmark - models ranked in order of performance using the average score (higher is better).
@@
GPT-4o-2024-08-06 0.4790 0.4990 0.5543 0.5318 0.5161
Llama-3.1-405B 0.3947 0.3974 0.4663 0.4883 0.4367
Qwen2.5-72B 0.3815 0.4370 0.4357 0.4410 0.4238
Qwen2.5-32B 0.3488 0.4052 0.4151 0.4481 0.4043
GPT-4o-mini-2024-07-18 0.3601 0.3327 0.4299 0.3659 0.3722
Llama-3.1-70B 0.3039 0.3141 0.3547 0.4253 0.3495
Gemma-2-27b 0.2963 0.3116 0.3675 0.3430 0.3296
Qwen2.5-14B 0.2289 0.2614 0.2786 0.3620 0.2827
Phi-3-medium 0.1745 0.1607 0.2154 0.2932 0.2110
Gemma-2-9b 0.1384 0.1269 0.1778 0.3092 0.1881
Qwen2.5-7B 0.1433 0.1311 0.1695 0.2849 0.1822
Llama-3.1-8B 0.1341 0.1582 0.1559 0.2637 0.1780
Phi-3-small 0.1168 0.1009 0.1531 0.1970 0.1419
Mistral-7B 0.0378 0.0417 0.0531 0.1789 0.0779
We report the strict scores of the medium, large and frontier models on the Lite Benchmark in Table 4. Unsurprisingly, GPT4o model performs the best on our benchmark data while large and medium-scale models like Llama-3.1 405B, Qwen2.5 72B, and, Qwen2.5 32B models appear to be better than other openly available models including Llama-3.1-70B-instruct and the Gemma family of models. We also include the results on the full benchmark in Appendix Table 5. We note that the ranking of models is largely consistent and that small models are much weaker than larger models.
@@ 5 Discussion & Conclusion
In this work, we demonstrated how modern LLMs fail to follow simple instructions. We took a novel approach to studying instruction-following by grounding instructions on existing knowledge tasks. Our approach has the advantage of being easily extendable for new instruction types and domains, while also enabling LLM-free evaluations with some degree of automated error analysis. We demonstrated that not only do models fail to follow simple instructions (e.g.) printing the answer text instead of the label, but their performance drops further when compound but simple, instructions are included. Even when instructions that should have no effect on the knowledge-tasks are used, models at all scales report a drop in performance, though the extent of deterioration varies. As models are increasingly being viewed as agents and assistants, it is crucial that models have better guarantees of following user instructions. As our benchmark demonstrates, there is a lot of scope for improvement and we hope the community finds it helpful in improving the current state-of-the-art.
Lastly, before concluding, we would like to re-emphasize the choice of the strict measure to study performance - if instructions specify how the task is to be completed then models should not add extraneous text, respond by rephrasing the question as part of the final response, make copying errors, etc. The nature of errors made by models as reflected in the difference between loose and strict scores, the automated error analysis sets and the large amount of unclassified errors highlights that instruction-tuning of LLMs requires special focus on instruction-following. We note that most existing instruction-tuning mixtures (eg: TuluV2 (Ivison et al., 2023)) focus on task-completion as opposed to instruction-following though recently some tuning mixtures have begun to include some tasks for instruction-following (Wang et al., 2024b). Datasets and methods such as those relying on the use of structured instruction prompts (Mishra et al., 2023) may also help aid instruction-following.
@@ Appendix A Appendix
We list all instructions with example input, ground truth, and expected instruction output in A.1. We the report results on the Full Benchmark in A.2. The detailed statistics of Full ad Lite Benchmark are presented in A.3. Section A.4 presents the comparison between model’s performance on print correct answer and print correct answer labels tasks on the Lite Benchmark. Section A.5 presents performance of different models for each instruction category in comparison with its corresponding performance on $print\_correct\_answer$ (PCA).
A.1 Prompts
@@
A.2 Full Benchmark LeaderBoard
We report strict scores of the small, medium, and, large models on the Full Benchmark in Table 5. We observe that Qwen2.5 72B and 32B variants outperforms all other models. Llama-3.1 70B is ranked third with a significant gap between the second best model Qwen2.5 32B. There is a significant drop in performance as the parameter size decreases.
Table 5: Performance of the Small, Medium, and Large Models on our Full Benchmark - models ranked in order of performance using the average score (higher is better).
Models
$\mu_{EM}$
IC
Score
KTS
Score
$\mu_{EM}^{{}^{\prime}}$
Average
Score
Qwen2.5-72B 0.4157 0.4827 0.4343 0.4697 0.4506
Qwen2.5-32B 0.3822 0.4501 0.4108 0.4846 0.4319
Llama-3.1-70B 0.3389 0.3532 0.3582 0.4623 0.3781
Gemma-2-27b 0.3392 0.3589 0.3727 0.3947 0.3664
Qwen2.5-14B 0.2545 0.2803 0.2767 0.3932 0.3012
Phi-3-medium 0.2013 0.1835 0.2167 0.3406 0.2355
Gemma-2-9b 0.1633 0.1583 0.1850 0.3409 0.2119
Qwen2.5-7B 0.1658 0.1515 0.1753 0.3366 0.2073
Llama-3.1-8B 0.1460 0.1741 0.1573 0.2927 0.1925
Phi-3-small 0.1352 0.1171 0.1568 0.2129 0.1555
Phi-3.5-mini 0.0883 0.0967 0.0932 0.1397 0.1045
Llama-3.2-3B 0.0704 0.0571 0.0780 0.1632 0.0922
Mistral-7B 0.0458 0.0541 0.0543 0.2093 0.0909
Qwen2.5-3B 0.0663 0.0704 0.0725 0.1090 0.0796
Qwen2.5-1.5B 0.0382 0.0347 0.0436 0.1223 0.0597
Llama-3.2-1B 0.0039 0.0037 0.0041 0.0184 0.0075
A.3 Additional Benchmark Statistics
@@
A.4 Printing the correct answer
@@
A.5 Knowledge-Task Characteristics and Instruction-Following
@@
Figure 7: Small Scale Models: Performance variation (strict and loose) of exact match scores for the Numeric Manipulation instruction category compared to its corresponding performance on $print\_correct\_answer$ (PCA).
@@
Figure 8: Large-Scale Models: Performance variation (strict and loose) of exact match scores for the Numeric Manipulation instruction category compared to its corresponding performance on $print\_correct\_answer$ (PCA).
Figure 9: Small-Scale Models: Performance variation (strict and loose) of exact match scores for the String Manipulation instruction category compared to its corresponding performance on $print\_correct\_answer$ (PCA).
Figure 10: Medium-Scale Models: Performance variation (strict and loose) of exact match scores for the String Manipulation instruction category compared to its corresponding performance on $print\_correct\_answer$ (PCA).
@@
Figure 11: Large-Scale Models: Performance variation (strict and loose) of exact match scores for the String Manipulation instruction category compared to its corresponding performance on $print\_correct\_answer$ (PCA).
@@
Figure 12: Frontier Models: Performance variation (strict and loose) of exact match scores for the String Manipulation instruction category compared to its corresponding performance on $print\_correct\_answer$ (PCA).
@@
Figure 13: Small-Scale Models: Performance variation (strict and loose) of exact match scores for the Format Correct Answer instruction category compared to its corresponding performance on $print\_correct\_answer$ (PCA).
@@
Figure 14: Medium-Scale Models: Performance variation (strict and loose) of exact match scores for the Format Correct Answer instruction category compared to its corresponding performance on $print\_correct\_answer$ (PCA).
@@
Figure 15: Large-Scale Models: Performance variation (strict and loose) of exact match scores for the Format Correct Answer instruction category compared to its corresponding performance on $print\_correct\_answer$ (PCA).
@@
Figure 16: Frontier-Scale Models: Performance variation (strict and loose) of exact match scores for the Format Correct Answer instruction category compared to its corresponding performance on $print\_correct\_answer$ (PCA).
@@
Figure 17: Small-Scale Models: Performance variation (strict and loose) of exact match scores for the Operations on List instruction category compared to its corresponding performance on $print\_correct\_answer$ (PCA).
@@
Figure 18: Medium-Scale Models: Performance variation (strict and loose) of exact match scores for the Operations on List instruction category compared to its corresponding performance on $print\_correct\_answer$ (PCA).
@@
Figure 19: Large-Scale Models: Performance variation (strict and loose) of exact match scores for the Operations on List instruction category compared to its corresponding performance on $print\_correct\_answer$ (PCA).
@@
Figure 20: Frontier Models: Performance variation (strict and loose) of exact match scores for the Operations on List instruction category compared to its corresponding performance on $print\_correct\_answer$ (PCA).
@@
Figure 21: Small-Scale Models: Performance variation (strict and loose) of exact match scores for the Operations on List (Conditional) instruction category compared to its corresponding performance on $print\_correct\_answer$ (PCA).
@@
Figure 22: Medium-Scale Models: Performance variation (strict and loose) of exact match scores for the Operations on List (Conditional) instruction category compared to its corresponding performance on $print\_correct\_answer$ (PCA).
@@
Figure 23: Large-Scale Models: Performance variation (strict and loose) of exact match scores for the Operations on List (Conditional) instruction category compared to its corresponding performance on $print\_correct\_answer$ (PCA).
@@
Figure 24: Frontier Models: Performance variation (strict and loose) of exact match scores for the Operations on List (Conditional) instruction category compared to its corresponding performance on $print\_correct\_answer$ (PCA).
A.6 Instruction Specific Results

### v2

@@ 4.2 Results
Figure 3: Impact of distractor instructions on exact match performance across tasks and instructions, compared to its corresponding $print\_correct\_answer$ performance. A drop indicates the model getting distracted by an inapplicable instruction. Results reported on Lite Benchmark.
(a) Llama Family of Models
(b) Qwen Family of Models
Figure 4: Classification of errors for the Llama and Qwen family of models.
@@ 4.2.1 Printing the correct answer
As mentioned earlier in Section 1 (Figure 1), we observe a significant drop ($\sim 20\%$ on average) in knowledge-task performance when instructing the model to respond with the text associated with the answer instead of its label. Some common issues we observed include models outright ignoring the instruction and continuing to generate labels, or generating only Chain-of-Thought reasoning without a final answer, etc, reflected by the knowledge and instruction following errors in Figure 2.
@@ 4.2.2 Answer-conditioned instruction performance
Table 3 reports the performance of all models on the print correct answer label (PCA Label) task which serves as the reference score for the knowledge/reasoning task.
Effect of Size Models: Starting with the Qwen 2.5 family of models it can be seen that in medium and large sized models (>14B parameter models) have comparable performance on the baseline answering task (PCA label). The drop for the large models (> 32B parameters) on the PCA task is relatively small as well. However, the drop in scores when combined with instructions result is much larger, and a trend is observed that larger models still do better than their smaller siblings. This holds true in all model families (Llama, Gemma and Phi) - even if there is a negligible difference in the PCA label performance of a model, the drop is less severe for a larger sibling from the same family. This result demonstrates that even if models perform comparably on the PCA label task, the large model size can benefit tasks which require knowledge and instruction composition.
Performance of Frontier Models: From Table 3 we find that frontier models also suffer from significant drops in performance. Notably, GPT4-o and Llama 3.1-405B Instruct have comparable performance on the knowledge task (as seen on PCA Label) but Llama’s performance deteriorates by a larger degree when combined with instructions.
4.2.3 Effect of distractor instructions
@@
Our dataset also includes instructions that apply only when certain properties of a knowledge-task answer are fulfilled. For instance, instructions for incrementing the correct answer by one if numeric, formatting numeric values, and printing any numeric answers in words, do not apply on tasks with textual answers. They serve as distractors, and we expect model performance to be unaffected since these instructions are not applicable and do not alter the original knowledge-task answer. From Figure 3, we observe that there is a 5-20% drop in small, medium, large, and frontier scale models.
In the appendix Figures 6, 7, 8, we report details of how different model families (Llama, Qwen, Phi) are affected by distractors, at different scales. In summary, we find that all models are distracted by instructions that require reversing and casing text,555partially applicable (can only reverse non-single digit numeric answers) or completely inapplicable reporting a drop of 55-75%. Even frontier models (Appendix Figure 9) show a drop of 35%. However, for the other distractor instructions, we find that Llama models are relatively robust, showing a drop of $5-10\%$, while the Phi and Qwen family of models are more susceptible, where we observe drops of 20-30%. Frontier models surprisingly also show a drop of 10-15%. While model failures in the presence of distractors have been studied before (Shi et al., 2023; Feng et al., 2024), to the best of our knowledge this is the first work to study them in an instruction-following setting.
4.2.4 Automated Error Classification
Figure 4 shows the error analysis for two different model families - Llama and Qwen2.5.666We include error analysis plots for other model families in the appendix. We report the average exact match score as well as the errors.
Errors and Model size: As can be seen in both Llama and Qwen2.5 model families, as the model size increases, the exact match accuracy increases and the errors reduce. The unclassified errors pertain to model outputs that are harder to classify and indicate a more severe model failure. As, the scale increases, such failures tend to reduce.
Instruction Categories and Errors: We note from Figure 4, the Llama and Qwen2.5 models have different strengths. For instance, in String Manipulation class of instructions, at very small scale the Llama 3.2-1B model and the Qwen2.5-1B model have similar exact match scores. However, in the case of the Llama model nearly 48% of all instances have errors that are unclassified (indicating severe failure) while that is not the case for the Qwen2.5 equivalent. Further, Qwen2.5-1B has a larger class of instruction-following errors as compared to knowledge errors; this suggests that the drop in performance for Qwen2.5-1B reported in Table 3 is indeed due to instruction-following errors made by the model.
In addition, we observe that the Qwen2.5 family of models in almost always better than the Llama family of models when it comes to numeric manipulation instructions while the Llama family of models has a slightly higher performance on string manipulation tasks (Appendix Figures 40 and 40). A similar trend is observed on the unclassified set where Llama’s instruction-following failures appear to be more severe. The Phi family of models (Appendix Figure 34) appears to perform worse than both the Qwen and Llama family of models. Frontier models (Appendix Figure 38) demonstrate lower knowledge errors as expected, but still make many instruction following errors.
@@ 4.3 Benchmark
GPT-4o-2024-08-06 0.5065 0.5174 0.5874 0.6889 0.575
Llama-3.1-405B 0.4617 0.4888 0.5351 0.6387 0.5311
Qwen2.5-72B 0.4348 0.5035 0.493 0.5768 0.502
Qwen2.5-32B 0.409 0.4751 0.4755 0.5873 0.4867
Llama-3.1-70B 0.3708 0.4138 0.4319 0.5645 0.4453
GPT-4o-mini 0.394 0.4029 0.4689 0.4609 0.4317
Gemma-2-27B 0.3497 0.3972 0.4194 0.4505 0.4042
Qwen2.5-14B 0.2764 0.3523 0.3272 0.4084 0.3411
Phi-3-medium 0.2518 0.2869 0.3054 0.4238 0.317
Gemma-2-9B 0.2381 0.2828 0.292 0.4428 0.3139
Qwen2.5-7B 0.1944 0.2513 0.2275 0.3411 0.2536
Llama-3.1-8B 0.174 0.2203 0.2048 0.3513 0.2376
Phi-3-small 0.1555 0.1809 0.1921 0.3027 0.2078
Mistral-7B 0.0577 0.0808 0.0768 0.205 0.1051
Table 4: Performance of the Medium, Large and Frontier Models on our Lite Benchmark - ranked in order of performance using the average score (higher is better).
We report the exact match scores of the medium, large and frontier models on the Lite Benchmark in Table 4. Unsurprisingly, the GPT4o model performs the best on our benchmark data while large and medium-scale models like Llama-3.1 405B, Qwen2.5 72B, and, Qwen2.5 32B based models appear to be better than other openly available models including Llama-3.1-70B-instruct and the Gemma family of models. We also include the results on the full benchmark in Appendix Table 5. We note that the ranking of models is largely consistent and that small models are much weaker than larger models.
@@ 5 Discussion & Conclusion
In this work, we took a novel approach to studying instruction-following by grounding instructions on existing knowledge tasks. Our approach has the advantage of being easily extendable for new instruction types and domains, while also enabling LLM-free evaluations with some degree of automated error analysis. We demonstrated that not only do models fail to follow simple instructions (e.g.) printing the answer text instead of the label, but their performance drops further when compound but simple, instructions are included. Even when instructions that should have no effect on the knowledge-tasks are used, models at all scales report a drop in performance, though the extent of deterioration varies. As models are increasingly being viewed as agents and assistants, it is crucial that models have better guarantees of following user instructions. As our work demonstrates, there is a lot of scope for improvement and we hope the community finds it helpful in improving the current state-of-the-art.
6 Limitations
One of the primary limitations of our approach is that if a model is not skilled in the underlying knowledge/reasoning task, the study of instruction-following when conditioned on the model answer being correct isn’t as helpful. Further, it is possible an LLM not considered in our set could generate responses in a style that is very incompatible with the assumptions made by our evaluation scripts. For example, some models could generate custom chain-of-thought markers which may require further adaptation of output processing and evaluation - we encourage users of our benchmark to inspect model outputs and update scripts if necessary.
@@ Appendix A Appendix
We begin by including details of the knowledge tasks used. We then describe our data creation process and how we automatically classify errors in section 3.4. We list all instructions with example input, ground truth, and expected instruction output in B.4. We report results on the Full Benchmark in Table 5. The detailed statistics of Full and Lite Benchmark are presented in B.5. Section C.1 presents the comparison between model’s performance on print correct answer and print correct answer labels tasks on the Lite Benchmark. Section C.3 presents performance of different models for each instruction category in comparison with its corresponding performance on $print\_correct\_answer$ (PCA).
Models
$\mu_{EM}$
IC
Score
KTS
Score
$\mu_{EM}^{{}^{\prime}}$
Average
Score
Qwen2.5-72B 0.4488 0.5077 0.4708 0.6218 0.5123
Qwen2.5-32B 0.419 0.4736 0.4519 0.6351 0.4949
Llama-3.1-70B 0.3697 0.3735 0.3925 0.6109 0.4366
Gemma-2-27B 0.3622 0.3783 0.3984 0.5177 0.4142
Qwen2.5-14B 0.2819 0.3521 0.305 0.4443 0.3458
Phi-3-medium 0.2589 0.2632 0.2799 0.4897 0.3229
Gemma-2-9B 0.2417 0.2701 0.2688 0.484 0.3162
Qwen2.5-7B 0.1921 0.2393 0.2 0.4061 0.2594
Llama-3.1-8B 0.1646 0.1917 0.1773 0.3907 0.2311
Phi-3-small 0.1472 0.1474 0.1686 0.3376 0.2002
Qwen2.5-3B 0.1277 0.1341 0.1386 0.3021 0.1756
Llama-3.2-3B 0.0946 0.0874 0.1021 0.2395 0.1309
Phi-3.5-mini 0.0966 0.1179 0.1014 0.2044 0.1301
Mistral-7B 0.0484 0.059 0.057 0.2451 0.1024
Qwen2.5-1.5B 0.0382 0.0346 0.0435 0.1461 0.0656
Llama-3.2-1B 0.0153 0.012 0.0176 0.0897 0.0337
Table 5: Performance of the Medium, Large and Frontier Models on our Full Benchmark - models ranked in order of performance using the average score (higher is better).
Appendix B Additional details - Dataset creation
B.1 Design Principles
We develop our instructions keeping the following design principles in mind: (i) We would like instructions to be unambiguous and be presented in a way that can be communicated clearly - if humans cannot follow the instructions and agree on the same output, LLMs should and likely would not be able to. (ii) We would like them to be easy to follow and not require complex reasoning abilities to follow so that models at all scales have a fair chance of success, (iii) The instructions need to have deterministic outputs that use the original answers of the knowledge-task or the candidate space of answers, or both, so that they can be evaluated easily with instruction specific scorers. (iv) We would like our benchmark to be based on a diverse mix of knowledge tasks, and be easily extensible to new ones.
B.2 Instruction Creation
To create each instruction, the authors iteratively refined them until all the authors had complete agreement in the output when they followed them manually. Examples of aspects of iterative improvement include - explicitly making clear what is not to be included in the output, how the output is to be presented, etc. We then asked $2$ computer science researchers to follow and generate the output for $75$ instructions across all our instruction types and datasets. We found that both the researchers were able to follow our instructions successfully and generated the same response for $93.33\%$ of the instances. The first annotator generated the correct response for $98.67\%$ of the instances, while the second annotator for $94.67\%$ of the instances. Upon analyzing their responses, we found the only instruction-following error was rounding off the decimal number when truncating to two decimal places. We also found very few human errors in the annotator’s response, specifically for instructions like $reverse\_correct\_answer\_alternate\_case$ on datasets with long output text such as PIQA.
B.3 Knowledge and Reasoning Tasks
(i) MMLUPro (Wang et al., 2024): MMLUPro extends the MMLU dataset to make it more challenging by a) increasing the number of options from four to ten and b) increasing problem difficulty by focusing on more reasoning oriented problems. We consider all $14$ subjects in the MMLUPro benchmark. We cap the maximum number of samples for each subject to be $150$ samples.
We also select a few common-sense and reasoning datasets:
B.4 Prompts
@@
B.5 Additional Benchmark Statistics
@@
Appendix C Additional Results
C.1 Printing the correct answer
@@
C.2 Influence of Distractors
Figures 6, 7 and 8 show the performance of the Llama, Qwen2.5 and Phi family of models in the presence of distractor instructions. Figure 9 shows the performance of frontier models.
Figure 6: Distractors on Llama Family of Models
Figure 7: Distractors on Qwen Family of Models
Figure 8: Distractors on Phi Family of Models
Figure 9: Distractors on Frontier Models
C.3 Knowledge-Task Characteristics and Instruction-Following
@@
Numeric Manipulation Instructions: Figures 10, 11 and 12 show the performance of small, medium and large models on the numeric manipulation category of instructions. These instructions only apply on the MathQA and MMLU datasets.
String Manipulation Instructions: Figures 14, 15, 16 and 17 show the performance of small, medium, large and frontier models on the string manipulation category of instructions. These instructions only apply on all datasets. Frontier models shown on data from Lite Benchmark.
Format Correct Answer Instructions: Figures 18, 19, 20 and 21 show the performance of small, medium, large and frontier models on the instructions that format the correct answer. These instructions only apply on all datasets. Frontier models shown on data from Lite Benchmark.
Operations on List: Figures 22, 23, 24 and 25 show the performance of small, medium, large and frontier models on the instructions operate on lists. These instructions only apply on all datasets. Frontier models shown on data from Lite Benchmark.
Operations on List (Conditional): Figures 26, 27, 28 and 29 show the performance of small, medium, large and frontier models on the instructions operate on lists. These instructions only apply on all datasets. Frontier models shown on data from Lite Benchmark.
@@
Figure 10: Small Scale Models: Performance variation of exact match scores for the Numeric Manipulation instruction category compared to its corresponding performance on $print\_correct\_answer$ (PCA).
@@
Figure 11: Medium Scale Models: Performance variation of exact match scores for the Numeric Manipulation instruction category compared to its corresponding performance on $print\_correct\_answer$ (PCA).
@@
Figure 12: Large-Scale Models: Performance variation of exact match scores for the Numeric Manipulation instruction category compared to its corresponding performance on $print\_correct\_answer$ (PCA).
@@
Figure 13: Frontier Models: Performance variation of exact match scores for the Numeric Manipulation instruction category compared to its corresponding performance on $print\_correct\_answer$ (PCA).
@@
Figure 14: Small-Scale Models: Performance variation of exact match scores for the String Manipulation instruction category compared to its corresponding performance on $print\_correct\_answer$ (PCA).
@@
Figure 15: Medium-Scale Models: Performance variation of exact match scores for the String Manipulation instruction category compared to its corresponding performance on $print\_correct\_answer$ (PCA).
@@
Figure 16: Large-Scale Models: Performance variation of exact match scores for the String Manipulation instruction category compared to its corresponding performance on $print\_correct\_answer$ (PCA).
@@
Figure 17: Frontier Models: Performance variation of exact match scores for the String Manipulation instruction category compared to its corresponding performance on $print\_correct\_answer$ (PCA).
@@
Figure 18: Small-Scale Models: Performance variation of exact match scores for the Format Correct Answer instruction category compared to its corresponding performance on $print\_correct\_answer$ (PCA).
@@
Figure 19: Medium-Scale Models: Performance variation of exact match scores for the Format Correct Answer instruction category compared to its corresponding performance on $print\_correct\_answer$ (PCA).
@@
Figure 20: Large-Scale Models: Performance variation of exact match scores for the Format Correct Answer instruction category compared to its corresponding performance on $print\_correct\_answer$ (PCA).
@@
Figure 21: Frontier-Scale Models: Performance variation of exact match scores for the Format Correct Answer instruction category compared to its corresponding performance on $print\_correct\_answer$ (PCA).
@@
Figure 22: Small-Scale Models: Performance variation of exact match scores for the Operations on List instruction category compared to its corresponding performance on $print\_correct\_answer$ (PCA).
@@
Figure 23: Medium-Scale Models: Performance variation of exact match scores for the Operations on List instruction category compared to its corresponding performance on $print\_correct\_answer$ (PCA).
@@
Figure 24: Large-Scale Models: Performance variation of exact match scores for the Operations on List instruction category compared to its corresponding performance on $print\_correct\_answer$ (PCA).
@@
Figure 25: Frontier Models: Performance variation of exact match scores for the Operations on List instruction category compared to its corresponding performance on $print\_correct\_answer$ (PCA).
Figure 26: Small-Scale Models: Performance variation of exact match scores for the Operations on List (Conditional) instruction category compared to its corresponding performance on $print\_correct\_answer$ (PCA).
Figure 27: Medium-Scale Models: Performance variation of exact match scores for the Operations on List (Conditional) instruction category compared to its corresponding performance on $print\_correct\_answer$ (PCA).
Figure 28: Large-Scale Models: Performance variation of exact match scores for the Operations on List (Conditional) instruction category compared to its corresponding performance on $print\_correct\_answer$ (PCA).
Figure 29: Frontier Models: Performance variation of exact match scores for the Operations on List (Conditional) instruction category compared to its corresponding performance on $print\_correct\_answer$ (PCA).
C.4 Instruction Specific Results